OpenStudy (anonymous):

What is the limit of y=x(9-x^2)^(1/2) as x approaches 3^- (AKA from the left)

OpenStudy (anonymous):

\[\lim_{x\to 3^{-}}x\sqrt{9-x^2}\]?

OpenStudy (anonymous):

Yes, that is it.

OpenStudy (anonymous):

since \(x<3\) you are in the domain of this function as \(x\to 3^-\) we have \(x^2\to 9\) and so \(9-x^2\to 0\) and likewise \(\sqrt{9-x^2}\to 0\)

OpenStudy (anonymous):

so your answer should be 0

OpenStudy (anonymous):

What was the process you used to yield that answer? Could you explain it? I'm not following is it is neither the dividing out method nor the rationalizing method.

OpenStudy (anonymous):

easier than that this is a product \(3\times 0=0\)

OpenStudy (anonymous):

So simply plugging in 3 would do the trick? No catch?